. The physician has ordered a heparin infusion of 12 Units/kg/hr. The client weighs 75 kg. The pharmacy sends an IV containing 50,000 units of heparin in 1000ml NS. The client will receive ___________ mL/min (Round to the first decimal place)
Answer #1

The physician has ordered a heparin infusion of 12 Units/kg/hr.

The client weighs 75 kg

= 75 x 12

= 900 /hr

= 15 units per minute

The pharmacy sends an IV containing 50,000 units of heparin in 1000ml NS

= 50,000/ 1000ml NS.

=50units/ml

= 15/50

= 0,3 ml/min